RFBG Commissioning with Adaptive Secondary - 23 March 2010 UT

Night 4: Investigating Z6 versus Elevation

Observer: JHill, JStorm (LBTO)
Telescope Operator: DHuerta
Software:
Instrument Support:: TShih
Telescope Support: JLittle
Adaptive Support:: MXompero (LBTO)

Summary

The night was mostly lost to heavy clouds. However we learned some interesting things in the brief time we were open.

We lost about an hour to the agw2-azcamg machine being turned off. See Issue 2633.

We confirmed that the excess astigmatism (Z6) that we see on the shell is a function of elevation. And Marco can see it on the shell positions as the telescope changes elevation. This means that the changing astigmatism is not caused by a changing shape of the reference body. Rather it must be caused by something in the control electronics. In the meantime, Marco has generated a lookup function to inject the compensating Z6 as a function of telescope elevation every 20 seconds.

I'm suspicious that there might be a varying coma term on the shell as well, but that is much more difficult to disentangle from the thermal and gravitational changes in collimation.

Andrew reports that the focus (Z4) measurements from his field data taken last night show essentially no tilt of the focal plane within the noise. This means that we are quite close to the null position for binodal astigmatism. Then he says: "Looking at the focus and astigmatism data we do have signs of binodal at the ~1 mm de-centre error level, (~20" on-sky M2 pointing) that we can address probably in one iteration after I take my calibration data."

Since we have a decent initial AGw probe transformation, and since we have decent optical alignment, and since we are able to do guiding and active optics with GCS......we (John, Andrew, Jesper) have concluded that we are ready to offer one or two half nights to Simone and friends to put the first starlight into the bigW unit for its initial sky checkout.

Near-term plan

It looks like the remainder of tonight (Monday) and Tuesday will be lost to weather. During Tuesday we will try to sort out where in the code the signs of the fixed field coma are being reversed (rotated). The priority for Wednesday night will be to collect more transformation data as that is Jesper's last night on the mountain.

I estimate that we are 30% complete on the AGW2 Commissioning activities (not including the Verification activities that follow). However, the majority of the remaining tasks require good seeing (refinement of transforms, refinement of binodal astigmatism, etc), so it is hard to say how many calendar nights remain. We probably need 4 good seeing nights.

Preparations

New drive module in RFBG rotator. (Experience during the night suggests that this has cured the rotator warnings that were causing problems.)

New Patch of PCS

Details

Jesper has trouble locating the cubes from last night. Because of the EOF problem, they get put in /newdata/indir/warning/[date] rather than in /Repository/[date]

01:45 $$$ restart of M3 UMAC to escape from selector stuck at +limit. Exact nature of problem unclear.

01:52 Restart GCSR per Torsten advice.

Open

Studying Pupil at High Elevation

01:59 pointto Pollux TelM="TRACK - no Agw unit selected

02:02 pointto Pollux TelM="TRACK - right authorize busy

02:03 pointto Pollux TelM="TRACK - Adjust IE from -48 to -53 (pointing model is good at EL=76.5).

02:05 pointto Pollux TelM="TRACK PAM-

M2 Zglobal=0.6 mm

Adjust M2 RX to -440 arcsec and M2 RY to 0 arcsec to remove coma. Adjust IE=-40 CA=-55 to compensate this repointing.

02:14 takeseq 0.2 40 (wrong date in cube name) irtc.20100323.021357.fits irtc.20100322_00070_cube.fit

02:15 takeseq 0.2 40 irtc.20100323.021451.fits irtc.20100323_00071_cube.fit

Z=2.6

takeseq 0.06 40 irtc.20100323.021901.fits cube 72

Cooling wasn't on for above cubes.

Focus is roughly at M2 Zglobal 1.73

02:23 pointto M5_0313 TelM="TRACK" PAM- no AGW unit selected

02:25 Stop and restart GCSR

02:26 pointto M5_0313 TelM="TRACK" PAM- - ptspiral - found star at IE=-50 CA=-124 at EL=19 IA from -314 to -390 to adjust Azimuth Zeropoint , and put CA back to -55 which was the high elevation value.

02:36 pointto M5_0313 TelM="ACTIVE" gs=0 no AGW unit selected.

Stop and restart GCSR

02:39 pointto M5_0313 TelM="ACTIVE" gs=0 - rejected change of guidestar list as GCS is currently switching or selecting an AGW unit.

02:41 Stop GCSR and then stopAgw startAgw

02:44 pointto M5_0313 TelM="ACTIVE" gs=0 - rejected change of guidestar list as GCS is currently switching or selecting an AGW unit.

stopAgw startAgw

02:50 pointto M5_0313 TelM="ACTIVE" gs=0 - rejected change of guidestar list as GCS is currently switching or selecting an AGW unit.

Both GCS and Torsten are out of commission. GCS has Issue 2633. Torsten is in bed with a high fever.

Studying Pupil at Low Elevation

03:03 Needed -20000 nm Z6 applied to M1 to make pupil look round. IA =-390 IE=-60 CA=-55

David reauthorizing (working on Issue 2633), so I have to restore my IA, IE, CA and reissue the preset.

03:12 pointto M5_0313 TelM="TRACK" PAM-

03:16 Marco applies +3 microns Z6 to M2 (This is the wrong sign.)

We changed that to -2.5 microns Z6 applied to M2 and the pupil looks OK (through very heavy clouds, so we didn't get a picture of it)

03:20 pointto M5_0313 TelM="ACTIVE" gs=0 - no AGW unit selected.

03:21 Close for heavy clouds.

Conclusion of Issue 2633

03:50 agw2-azcamg was not pingable. We found that it was powered off. Probably it got bumped during the M3 UMAC debugging happening directly adjacent to it this afternoon.

04:04 pointto M5_0313 TelM="ACTIVE" gs=0 - no AGW unit selected.

Stop GCSR, stopAGW, startAGW, John had to soft restart the agw2-azcamw machine. (This was because Norm had logged out of it a few minutes earlier while checking the password.)

04:16 Restart GCSR again.

Mar 23 04:15:53 tcs4 LBT_right_GCS: AGW selection succeeded.  side: R, Instrument: IRTC, FS: bentGregorianFront, initialized from file 'AIP_IRTC_R.cfg'
Mar 23 04:15:53 tcs4 LBT_right_GCS: successfully recovered from previous crash. AGW unit selected based on currently authorized instrument/FS
Mar 23 04:15:53 tcs4 LBT_right_GCS: current AGW unique ID is: 10.144.0.105:2

04:18 pointto M5_0313 TelM="ACTIVE" gs=0 - succeeded up to not finding a star on the inside of a dome.

Open

05:10 Open to a partly cloudy sky

05:17 pointto M5_0384 gs=0 TelM="ACQUIRE" PAM- EL=80 IE=-48 CA=-40

05:21 pointto M5_0384 gs=0 TelM="TRACK" PAM- Change RX from -440 to -520. IE to -36 384 doesn't have guide stars.

05:30 pointto M5_0358 gs=1 TelM="ACTIVE" - didn't find the guide star - clouds?

05:36 pointto M5_1301 gs=0 TelM="TRACK" - EL=35

05:40 pointto M5_1301 gs=0 TelM="TRACK" PAM- - CA=-42 RX=-400 IE=-55

Send to M1 -10000 nm Z6 (pupil is round to your eye)

05:47 takeseq 0.5 40 irtc.20100323.054721.fit cube 73

Clear C00 M1 forces

05:48 takeseq 0.5 40 irtc.20100323.054826.fits cube 74

05:49 Marco applies +1.5 microns Z6 to M2

05:50 takeseq 0.5 40 irtc.20100323.054954.fits cube 75

05:51 Marco applies +2.0 microns Z6 to M2 EL=36

05:52 takeseq 0.5 40 irtc.20100323.055131.fits cube 76

05:53 Slew to zenith, and then back to this target. (Z6 seems unchanged by eye) Marco says that he sees the astigmatism shape appear on the shell while the telescope is slewing. Thus, it is not a problem of the reference body. Marco's theory is that the astigmatism is created because the integrators on the actuators are disabled in one sector at the bottom of the mirror in order to stabilize the shell in the area where there are actuator problems.

05:57 takeseq 0.5 40 irtc.20100323.055730.fits cube 77

06:04 pointto BS9148 gs=0 TelM="ACQUIRE" PAM- Adjust IE=-66 CA=-36

06:05 Andrew calls to say that we certainly have coma reversed on the compensation of fixed field aberrations. Doug can correct that in the forthcoming snow storm.

06:09 Lose 8 minutes when John's X-server freezes on obs2. Fortunately he remembers the infamous Cntrl-Alt-Backspace command to kill the X-server and logout. (Thanks Stephen and Norm.)

Close

06:15 Close for heavy clouds. T=0.9 degC. Based on the satellite photo, this may be the last observing until Wednesday night.

Testing the RFBG rotator module closed dome.

06:45 pointto BS9152 gs=0 TelM="TRACK"

06:58 David stopped the rotator.

07:38 pointto BS9152 gs=0 TelM="TRACK" again

07:45 Marco has completed the first cut program which applies Z6 to the shell every 20 seconds as the telescope elevation changes. We are testing it closed-dome while the telescope is tracking during these rotator tests.

Snow

11:50 No accumulation yet, but the snow flurries have started.

-- JohnHill - 23 Mar 2010
Topic revision: r3 - 23 Mar 2010, JohnHill
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback